Daily Report
Back to Expedition
Exp.358 Daily Report

February

2.20, 2019
Day:137

Location: Site C0002
Weather: Fine but cloudy

Using wireline to retrieve core sample from the depth, drill pipe remain in the bore hole.

Today's photo:
Inner core barrel on the deck.

2.18, 2019
Day:135

Location: Site C0002
Weather: Fine but cloudy

Cutting the first core last night, and recovered on the deck early this morning.

Today's photo:
Onboard scientists looking into the first core.

2.17, 2019
Day:134

Location: Site C0002
Weather: Cloudy

Running coring BHA, although we can not recover the plate boundary fault core sample, we try to retrieve core sample from the accretionary prism.

Today's photo:
Deck crew preparing coring BHA.

2.14, 2019
Day:131

Location: Site C0002
Weather: Cloudy

Data and sample analysis and report writing is ongoing in the lab.

Today's photo:
Logging staff scientist checking LWD data.

2.13, 2019
Day:130

Location: Site C0002
Weather: Fine but cloudy

Onboard scientists are obligated to write expedition report. Not only cuttings analysis but also LWD data are analyzed.

Today's photo:
Logging scientist checking LWD data.

2.12, 2019
Day:129

Location: Site C0002
Weather: Fine but cloudy

Measurement and analysis of cuttings continues in the lab.

Today's photo:
Lab tech making thin section of cuttings.

2.11, 2019
Day:128

Location: Site C0002
Weather: Fine but cloudy

Information from Chikyu

Today's photo:
Micropaleontologist processing cuttings for radioralian observation.

2.8, 2019
Day:125

Location: Site C0002
Weather: Fine but cloudy

Hole condition is good, continue drilling.

Today's photo:
Real-time-monitoring team looking bore hole image data.

2.7, 2019
Day:124

Location: Site C0002
Weather: Rain

Continue drilling ahead with new BHA, LWD data as well as cuttings are coming onboard.

Today's photo:
Fresh cuttings at the shale shaker room.

2.6, 2019
Day:123

Location: Site C0002
Weather: Fine but cloudy

Continue LWD BHA at Hole C0002S, LWD data is coming onboard.

Today's photo:
Science Leaders watching real time LWD data at onboard lab.

2.5, 2019
Day:122

Location: Site C0002
Weather: Overcast

Running the next BHA with LWD tools into the hole.

Today's photo:
New drilling bit for the current BHA.

2.4, 2019
Day:121

Location: Site C0002
Weather: Partly cloudy

Conduct gyro survey and start kick-off the hole.

Photo from archive:
Beautiful sunset with supply boat.

2.3, 2019
Day:120

Location: Site C0002
Weather: Partly cloudy

Running the next kick-off BHA into the hole.

Today's photo:
Retrieved Tri-mill BHA at the rig floor. Good job!

2.2, 2019
Day:119

Location: Site C0002
Weather: Partly cloudy

Dressing the window has successfully finish, pull out the BHA from the hole.

Today's photo:
The moon, Mercury, and Venus all appearing in the dawn sky above the Chikyu's crane.

2.1, 2019
Day:118

Location: Site C0002
Weather: Partly cloudy

Set the whipstock and window milling was successfully finish.

Today's photo:
Scientists are on a morning walk on the helideck.


January

1.31, 2019
Day:117

Location: Site C0002
Weather: Partly cloudy

Running whipstock assembly with window mill for casing window cut.

Today's photo:
Scientists join the crowd on the rig floor as the whipstock assembly is rigged up and run in hole.

1.30, 2019
Day:116

Location: Site C0002
Weather: Clear

Drilling equipments maintenance and wireline survey inside the casing successfully done.

Today's photo:
Group photo of the current science party members.

1.29, 2019
Day:115

Location: Site C0002
Weather: Partly cloudy

Finish retrieving the current BHA. After maintenance of drilling equipments, wireline tool surveying inside the casing.

Photo from archive:
She brings us drilling equipments, food and fuel for us.

1.28, 2019
Day:114

Location: Site C0002
Weather: Cloudy

While retrieving the current BHA, BOP (Blow Out Preventer) pressure test has been conducted, confirmed working properly.

Today's photo:
Videographer filming scientists working in the lab.

1.27, 2019
Day:113

Location: Site C0002
Weather: Partly cloudy

Starting retrieve the current BHA.

Today's photo:
Scientists engaged in the Ping Pong tournament in their (limited) off time.

1.25, 2019
Day:111

Location: Site C0002
Weather: Partly cloudy

Retrieving the current BHA from the well is ongoing.

Today's photo:
Drilling crew working on assembling next wireline drill pipe clearing tool.

1.24, 2019
Day:110

Location: Site C0002
Weather: Partly cloudy

Current BHA retrieving plan has been made, and preparation started.

Today’s photo:
Video crew filming scientists rig tour on the top of the derrick.

1.23, 2019
Day:109

Location: Site C0002
Weather: Partly cloudy

Tel conference with shore base regarding future operation plan took place. Meanwhile, new science party members visit rig floor.

Photo from archive:
Science leader visiting the rig floor, you may not catch a bus here.

1.22, 2019
Day:108

Location: Site C0002
Weather: Partly cloudy

One of the science leader replaced today. She sailed on about 1.5 months. Otsukaresamadesita!

Today’s photo:
Drilling Engineer displays the sponge ball used during cementing operations.

1.21, 2019
Day:107

Location: Site C0002
Weather: Partly cloudy

Newly embarked science party members are start working on new measurements of cuttings.

Today’s photo:
Scientists work on preparing samples for moisture and density measurements.

1.18, 2019
Day:104

Location: Site C0002
Weather: Partly cloudy

Ten new science party members were embark today, onboard laboratories are become lively.

Today’s photo:
Science Leader taking a moment for silent contemplation during the sunset.

1.17, 2019
Day:103

Location: Site C0002
Weather: Partly cloudy

Checking the condition of the bottom of the casing, meanwhile slip and cut of the drilling line has been done.

Today’s photo:
Drilling crew working for the drilling line slip and cut.

1.16, 2019
Day:102

Location: Site C0002
Weather: Partly cloudy

The operations have hit a bit of problem with drilling out the cement, so we are still working on overcoming the issue.

Photo from archive:
Lead scientist meets ROV which normally underwater.

1.15, 2019
Day:101

Location: Site C0002
Weather: Cloudy

The casing expansion was successfully completed and the steps to drill out cement began.

Photo from archive:
Lead scientist seeing expandable casing.

1.14, 2019
Day:100

Location: Site C0002
Weather: Partly cloudy

Expandable casing was successfully set, although it was shallower than planned.

Today's photo:
Cuttings samples prepared for personal sample requests.

1.13, 2019
Day:99

Location: Site C0002
Weather: Partly cloudy

Expandable casing is running into the hole to covering troublesome layers.

Today's photo:
We had Kaisen-Chirashi for dinner tonight, looks pretty and tastes good!

1.12, 2019
Day:98

Location: Site C0002
Weather: Partly cloudy

Based on the data collected by LWD, some intervals of the well were enlarged.

Today's photo:
Scientists and OSI examine recovered underreamer.

1.11, 2019
Day:97

Location: Site C0002
Weather: Partly cloudy

After a day spent clearing and sweeping out the hole, the BHA is pulling out to the surface.

Today's photo:
Scientists examine cuttings sample using a microscope.

1.10, 2019
Day:96

Location: Site C0002
Weather: Partly cloudy

Running the reaming BHA to the hole, but it is difficult to go deeper.

Today's photo:
Roughneck on the rigfloor working on the pipehandler.

1.9, 2019
Day:95

Location: Site C0002
Weather: Partly cloudy

Stop enlarging the hole, recover the LWD/reamer BHA on the deck.

Today's photo:
The OSI and Engineers examine the damage on the recovered reamer.

1.8, 2019
Day:94

Location: Site C0002
Weather: Partly cloudy

Reaming upper section close to the casing window for make sure the hole is in ok condition.

Today's photo:
Some of the ingredients for drilling mud include these huge bags of salt - 1000 kg!

1.7, 2019
Day:93

Location: Site C0002
Weather: Partly cloudy

LWD-Reamer BHA drilling is ongoing for section 1.

Today's photo:
Science Leader practices drop the ball into the drill pipe on the rig floor to activate the underreamers.

1.6, 2019
Day:92

Location: Site C0002
Weather: Partrly cloudy

The new LWD BHA was made up and run into the rig floor, waiting for dolly maintenance on the top drive.

Today's photo:
As part of Sunday safety drill, a lifeboat swingout checks were conducted.

1.5, 2019
Day:91

Location: Site C0002
Weather: Partrly cloudy

Kick-off BHA layout on the rig floor, after BOP function test.

Today's photo:
Drilling crew checking retrieved mud motor assembly.

1.4, 2019
Day:90

Location: Site C0002
Weather: Partrly cloudy

Finish drilling with the kick-off BHA, while pulling out of the hole, conduct BOP function test.

Today's photo:
the roughnecks working to break the saver sub on the rigfloor.

1.3, 2019
Day:89

Location: Site C0002
Weather: Fine but cloudy

Drilling down to kick off BHA target depth, and clean up the hole.

Today's photo:
One of the science leader visits the moon pool.

1.2, 2019
Day:88

Location: Site C0002
Weather: Fine but cloudy

Drilling in progress, we will reach the target depth of this pilot hole soon.

Today's photo:
When drilling is going well, the rig floor is quieter.

1.1, 2019
Day:87

Location: Site C0002
Weather: Fine but cloudy

Drilling with the 2nd kick off BHA, Based on the survey data, the hole inclination gradually become vertical.

Today's photo:
Shale shaker where cuttings (sample) come up from the well.
December

12.31, 2018
Day:86

Location: Site C0002
Weather: Fine but cloudy

Happy New Year! People gathering on helideck to see the first sunrise, although it was cloudy in the horizon.Best wishes for 2019!

12.30, 2018
Day:85

Location: Site C0002
Weather: Fine but cloudy

Running into the hole with 2nd kick-off BHA with higher motor bent angle (1.15 - 1.50 degree).

Today's photo:
Slip and cut drill line.

12.29, 2018
Day:84

Location: Site C0002
Weather: Fine but cloudy

Recover 8-1/2” kick off BHA to the surface for changing drilling bit.

Today's photo:
Onboard science team (lithostratigraphy) discussing at the core lab.

12.28, 2018
Day:83

Location: Site C0002
Weather: Fine but cloudy

Continue drilling with the kick-off BHA. Six new scientists and EPM had embarked.

Today's photo:
Kick-off meeting of the science party at the lab conference room.

12.27, 2018
Day:82

Location: Site C0002
Weather: Fine but cloudy

Continuously working with the kick-off BHA, hole condition is not bad.

Today's photo:
A lab tech putting a cuttings sample into a brand-new powerful vacuum dryer.

12.26, 2018
Day:81

Location: Site C0002
Weather: Cloudy

Continuing drilling with the kick-off BHA.

Today's photo:
15 laboratory technicians are onboard, working for sampling and several measurements.

12.25, 2018
Day:80

Location: Site C0002
Weather: Fine but cloudy

Continue drilling with the kick-off BHA.

Today's photo:
Mud-gas scientists heading the lab container for hand-over.

12.24, 2018
Day:79

Location: Site C0002
Weather: Fine but cloudy

Continue drilling into fragile formation with the kick-off BHA.

Today's photo:
OSI explaining the current drilling plan to the science leaders.

12.23, 2018
Day:78

Location: Site C0002
Weather: Fine but cloudy

The kick-off BHA started drilling Hole 2R last night.

Today's photo:
A mud-gas scientist sampling formation gas in a tube in Mud-gas Monitoring Lab.

12.22, 2018
Day:77

Location: Site C0002
Weather: Fine but cloudy

Casing window dressing was successfully done, next BHA will be a kick-off assembly.

Today's photo:
washing the window mill assembly to check the condition.

12.21, 2018
Day:76

Location: Site C0002
Weather: Fine but cloudy

Casing window dressing operation continues, this is really important for the future casing set.

Today's photo:
One of the science leader changed, this is the ultra-short hand-over between two science leaders.

12.20, 2018
Day:75

Location: Site C0002
Weather: Fine but cloudy

Prior to drilling deeper, casing dressing BHA is running into the hole.

Today's photo:
Scientists having discussion at the front of composite cuttings in a core liner in the Core Lab.

12.19, 2018
Day:74

Location: Site C0002
Weather: Fine but cloudy

After milling operation with the 2nd BHA, a Formation Integrity Test (FIT) was conducted with good results.

Today's photo:
A meeting with the onboard scientists, lab officer, and the lab curator in the Co-Chief office discussing additional measurements.

12.18, 2018
Day:73

Location: Site C0002
Weather: Fine but cloudy

Continue casing window opening.

Today's photo:
Science Leader is excited to examine the reamer activation ball before it's dropped to activate the downhole reamers.

12.17, 2018
Day:72

Location: Site C0002
Weather: Fine but cloudy

Running whipstock assembly with window mill for casing window cut.

Today's photo:
OSI briefing new operation plan to onboard Science Leaders.

12.16, 2018
Day:71

Location: Site C0002
Weather: Fine but cloudy

Decision was made for new side-tracking due to the current hole condition.

Today's photo:
OSI carefully watching whipstock and tri-mil assembly.

12.15, 2018
Day:70

Location: Site C0002
Weather: Fine but cloudy

Drilling equipments maintenance are conducted, and preparation of the next drilling is ongoing.

Today's photo:
Two science leader visiting mud-gas monitoring lab container onboard.

12.14, 2018
Day:69

Location: Site C0002
Weather: Fine but cloudy

We face some difficulty with the current BHA, so pull out the hole and change the BHA.Meanwhile, sample curation are ongoing in the lab.

Today's photo:
Lab tech register the samples one by one to the database.

12.13, 2018
Day:68

Location: Site C0002
Weather: Fine but cloudy

Window 1 Science party were almost disembarked, and only science leaders and EPM are in the lab.

Today's photo:
Science Leader taking samples for a personal request.

12.12, 2018
Day:67

Location: Site C0002
Weather: Fine but cloudy

In the laboratory, the shipboard SLs themselves sorted and selected cuttings for personal requests.

Today's photo:
1st cuttings sampling party

12.11, 2018
Day:66

Location: Site C0002
Weather: Overcast

The first group of scientists disembarked today, only the Science Leaders & a few specialists remain aboard.

Today's photo:
Encouraging messages of support from the public arrive onboard from shore, & are displayed in the lab.

12.10, 2018
Day:65

Location: Site C0002
Weather: Cloudy

Running a new BHA & experiencing some delays in progress; drilling into a new formation might be the cause, but we need steady progress!

Today's photo:
Christmas tree decoration in the galley.

12.9, 2018
Day:64

Location: Site C0002
Weather: Fine but cloudy

Disembarkation schedule for the current science party is coming soon. Discussion of the scientific results are ongoing in the lab.

Today's photo:
A structure geologist showing features of cuttings.

12.8, 2018
Day:63

Location: Site C0002
Weather: Fine but cloudy

All the BHA was pulled out of the hole, changing to the new BHA.

Today's photo:
Drilling bit used for the previous BHA.

12.7, 2018
Day:62

Location: Site C0002
Weather: Overcast

Drill string pulling out from the hole for changing the BHA. Continue to drill with the new BHA.

Today's photo:
New EPM (left) and new science leader embarked yesterday.

12.6, 2018
Day:61

Location: Site C0002
Weather: Overcast

We use helicopter for transportation of scientistsand crew from Minami-Ise heli base for this long expedition.

Today's photo:
The helicopter waiting for boarding.

12.5, 2018
Day:60

Location: Site C0002
Weather: Cloudy

Good communication is the key of the success. Not only the daily morning tel conferences, occasional tel conference between Chikyu and onshore takes place even in the night time.

Today's photo:
Daily morning tel conference.

12.4, 2018
Day:59

Location: Site C0002
Weather: Cloudy

Keep drilling ahead, steadily increasing the penetration rate, but still less than planned.

Today's photo:
Onboard scientists celebrate the achievement of a new depth record in scientific ocean drilling.

12.3, 2018
Day:58

Location: Site C0002
Weather: Clear

Passed the old C2P depth record (3058.5 mbsf) last night, we are now in unknown territory.

Today's photo:
Operation Superintendent (OSI) discussing with a drilling crew.

12.2, 2018
Day:57

Location: Site C0002
Weather: Rainy

Currently 14 scientists are onboard, divided to two shifts (00:00 - 12:00 and 12:00 - 24:00) and continuously working in the lab. Everyday after 11:00, all the members gather and holding update meeting.

Today's photo:
Science party members (plus alfa) at the moment.

12.1, 2018
Day:56

Location: Site C0002
Weather: Clear

There is a library in the lab where scientists writing report and discuss. This library called “the professor Munk Memorial Library” after Prof. Walter Munk who proposed the mantle drilling.

Today's photo:
A scientist discuss in the library.


November

11.30, 2018
Day:55

Location: Site C0002
Weather: Partly cloudy

Revise the original plan of changing the BHA, decision was made to keep drilling to the TD of section 1 with the current BHA.

Today's photo:
A scientist delivers some updates on XRD data to the science party.

11.29, 2018
Day:54

Location: Site C0002
Weather: Partly cloudy

Slow but steady progress of drilling operation.

Today's photo:
Mechanics working in the core tech work shop.

11.28, 2018
Day:53

Location: Site C0002
Weather: Rain

Continue drilling with the shorter BHA but does not progress as we expected. But no rush, keep it steady.

Today's photo:
One of two hawks that make the derrick aboard Chikyu their home.

11.27, 2018
Day:52

Location: Site C0002
Weather: Clear

As an obligation of the science party member, onboard scientists need to write the expedition report. Current section 1 method chapter due this Friday.

Today's photo:
Supply boat “Akatsuki”, she brings drilling equipments, food, and fuel to us.

11.26, 2018
Day:51

Location: Site C0002
Weather: Partly cloudy

The current bottom of the hole is about 5,000 m from the ship, it took about 20 hours one way for the trip.

Today's photo:
Mid-day science meeting in the lab.

11.25, 2018
Day:50

Location: Site C0002
Weather: Partly cloudy

Somehow the LWD-BHA interfere with the casing window. Re-challenge with a shorter LWD-BHA.

Today's photo:
Rig tour for the onboard scientists.

11.24, 2018
Day:49

Location: Site C0002
Weather: Partly cloudy

Real time LWD data were collected. After tool recovery, downloaded memory data will send to the shore for further analysis.

Today's photo:
All the drilling operation takes place here, the driller’s house.

11.23, 2018
Day:48

Location: Site C0002
Weather: Partly cloudy

Ready for the samples/data from C0002Q well in the lab.

Today's photo:
Sunrise from the helideck with the morning star (Venus).

11.22, 2018
Day:47

Location: Site C0002
Weather: Partly cloudy

Conducting LWD shallow tests and functional test of drilling tools while running into the hole.

Today's photo:
Scientists listening the instructions on preparation techniques in the chem lab.

11.21, 2018
Day:46

Location: Site C0002
Weather: Cloudy

Start making up the next BHA that includes LWD (Logging-While-Drilling) tool. We obtain several information from the formation using this tool while drilling.

Today's photo:
Making up the LWD tool at the drill floor.

11.20, 2018
Day:45

Location: Site C0002
Weather: Partly Cloudy

10 scientists from 4 nations embarked. Start working in the lab area soon after the safety induction.

Today's photo:
Onboard scientists examining structure in cuttings.

11.19, 2018
Day:44

Location: Site C0002
Weather: Partly Cloudy

Drilling with the kick-off BHA has been finished, cleaning the hole.

Today's photo:
Lab curator demonstrating cuttings sieving.

11.18, 2018
Day:43

Location: Site C0002
Weather: Cloudy

Kick-off BHA start drilling from 2,890 m below sea floor.

Today's photo:
Kick-off BHA at the rig floor, No.2

11.17, 2018
Day:42

Location: Site C0002
Weather: Partly Cloudy

Start running kick-off BHA. Drill a new hole from the casing window.

Today's photo:
Kick-off BHA at the rig floor.

11.16, 2018
Day:41

Location: Site C0002
Weather: Partly Cloudy

Casing milling BHA was retrieved on the rig floor. After replacing some parts of the traveling block dolly, we will run a kick off BHA.

Today's photo:
Engineers are checking the dressing bit for wear and tear.

11.15, 2018
Day:40

Location: Site C0002
Weather: Partly Cloudy

Finally casing window milling operation was done. Meanwhile, holding a meeting in the lab regarding mud gas sampling.

Today's photo:
Worn out milling bit.

11.14, 2018
Day:39

Location: Site C0002
Weather: Partly Cloudy

Casing milling operation continues.

Today's photo:
New BHA for milling operation.

11.13, 2018
Day:38

Location: Site C0002
Weather: Cloudy

Still working on dressing the casing window in preparation for eventual expandable casing running.

Today's photo:
Serious eyes of an engineer on the rig floor.

11.12, 2018
Day:37

Location: Site C0002
Weather: Rain

After milling the casing, Formation Integrity Test (FIT) was conducted and the test results was good!

Today's photo:
Smile! Tentative seal for the riser pipe at the rig floor.

11.11, 2018
Day:36

Location: Site C0002
Weather: Partly Cloudy

Casing milling operation is almost done with the second milling bit.

Today's photo:
The milling BHA bit shows the wear and tear from trying to cut out a hole in the 11-3/4-inch casing.

11.10, 2018
Day:35

Location: Site C0002
Weather: Partly Cloudy

After one day delay, OSI and LSS were replaced. Meanwhile, casing milling operation continues.

Today's photo:
Cuttings collection team in the Shale Shaker Room.

11.9, 2018
Day:34

Location: Site C0002
Weather: Partly Cloudy

Continue casing milling operation. Scheduled tel conference among science leaders were canceled due to bad weather.

Today's photo:
Approaching rain front off Chikyu’s helideck.

11.8, 2018
Day:33

Location: Site C0002
Weather: Partly Cloudy

Running whipstock second time is now on going, things goes well, we will start drilling a new hole (Hole Q) soon.

Today's photo:
Drilling Sequences

11.7, 2018
Day:32

Location: Site C0002
Weather: Partly Cloudy

Dry run walkthrough of cuttings flow in lab is conducted, the results is reflected to the cuttings flow.

Today's photo:
The ROV, working at the sea bottom.

11.6, 2018
Day:31

Location: Site C0002
Weather: Rain/PartlyCloudy

One of the science leader and Expedition Project Manager (EPM) replaced. In addition, two more scientists embarked today.

Today's photo:
Drilling crew working in night shift.

11.5, 2018
Day:30

Location: Site C0002
Weather: Overcast

The first meeting of RTG Team (Real Time Geomechanics Team) was held for recommendation of kick-off angle to the formation with both onboard and onshore team members. Then whipstock and window cut mill assembly is running into the hole.

Today's photo:
Window cut mil.

11.4, 2018
Day:29

Location: Site C0002
Weather: Fine but cloudy

Confirmed the condition of cement outside the casing, and it was good. Prepare for kick-off operation.

Today's photo:
A scientist visiting wireline logging control van.

11.3, 2018
Day:28

Location: Site C0002
Weather: Cloudy

Scraping inside of two previously set casing was completed. Next step is run a logging tool to check cement bond outside the casing.

Today's photo:
wireline tool for cement bond check.

11.2, 2018
Day:27

Location: Site C0002
Weather: Cloudy

Inside casing cleaning job was done using a scraper BHA (Bottom-Hole-Assembly). Tel conference for all the science party members in the morning. Continuous cuttings handling preparation in the lab area.

Today's photo:
cuttings handling instructions.

11.1, 2018
Day:26

Location: Site C0002
Weather: Fine but cloudy

Three shoe bond tests has been conducted at 11-3/4 inch liner casing shoe. Next step is cleaning both 13-3/8 and 11-3/4 inch casings.

Today's photo:
drill strings standing at pipe racker under the derrick.


October

10.31, 2018
Day:25

Location: Site C0002
Weather: Fine but cloudy

Continue drill-out cement and cleaning 11-3/4 inch liner casing. Preparation for the cuttings handling and measurement continues in the lab area.

Today's photo:
set "rotary slip" to hold the drill string.

10.30, 2018
Day:24

Location: Site C0002
Weather: Fine but cloudy

Finish drill out shallower cement plug in 13-3/8 inch casing and run 10-5/8 inch drill-out-cement BHA (Bottom-Hole-Assembly) for the deeper part of 11-3/4 inch liner casing.

Today's photo:
checking engine operation status at the engine control room

10.29, 2018
Day:23

Location: Site C0002
Weather: Fine but cloudy

Drilling out the cement plug inside 13-3/8" casing and pumping new mud into the well was completed. Now making and running the next drill-out-cement BHA in hole.

Today's photo:
12-1/2 inch drill-out-cement bit.

10.28, 2018
Day:22

Location: Site C0002
Weather: Fine but cloudy

Drill out cement plug and cleaning inside casing continues. In the lab area, continue preparation for cuttings handling tools and bottles.

Today's photo:
Loading materials (should include fresh food!) from a supply boat Akatsuki

10.27, 2018
Day:21

Location: Site C0002
Weather: Cloudy

Conduct pressure test on 13-3/8” casing, and drill out cement plug continues. Live broadcast from onboard Core Lab for JAMSTEC Yokohama open house event to introduce NanTroSEIZE project for public audience. We met many good questions from the participants; for example, “why does earthquake happen” from a 4-year-old kid, and Science Leaders Q-san and Masa-san replied and explained to the questions.

Today's photo:
Live broadcasting with DV Chikyu at JAMSTEC Yokohama Institute

10.26, 2018
Day:20

Location: Site C0002
Weather: Rain

Run Drill Out Cement Bore Hole Assembly (DOC BHA), wash down to tab cement top, and replace drilling mud. Meanwhile, hold a Science Leaders meeting to discuss scientists embarkation schedule, sample request approval, and science party meeting.

Today's photo:
connecting drill-out-cement bit.

10.25, 2018
Day:19

Location: Site C0002
Weather: Fine but cloudy

Install hydraulic power swivel on traveling block, make up and run wear bushing running retrieving tool (WBRRT), conduct pressure test, pull WBRRT out of hole. Continue preparation for cuttings handling tools and bottles in the lab area.

Today's photo:
working on the hydraulic power swivel before running.

10.24, 2018
Day:18

Location: Site C0002
Weather: Overcast

Conduct riser rotation test, install Upper Flex Joint and Diverter, rig down riser handling equipment. In the lab area, continue preparation for cuttings handling tools and bottles.

Today's photo:
Riser Tensioner hangs riser joints.

10.23, 2018
Day:17

Location: Site C0002
Weather: Overcast

Secure multiplexer cables, latch tensioner ring on telescopic joint, land BOP on wellhead, conduct pressure test on wellhead connector. Science Leader Prof. Masa Kinoshita embarked in the afternoon and had a short meeting for updates. There was crew change for the half of lab techs, and they have also update/handover meeting.

Today's photo:
just before landing

10.22, 2018
Day:16

Location: Site C0002
Weather: Overcast

Continue running BOP/riser, install Gooseneck and Moon-pool hose. Conduct pressure tests on Auxiliary Lines. Continue preparation for cuttings handling tools and working spaces.

Today's photo:
connecting Moon-pool hose to riser joint.

10.21, 2018
Day:15

Location: Site C0002
Weather: Fine but cloudy

BOP/Riser running continues smoothly, and close to the well head. Meanwhile cuttings handling dry-run took place in the lab area.

Today's photo:
Riser pipe handling onboard.

10.20, 2018
Day:14

Location: Site C0002
Weather: Fine but cloudy

BOP/Riser running continues smoothly. Meanwhile reviewing layout and preparation for cuttings in the lab area.

Today's photo:
pick up riser by a crane.

10.19, 2018
Day:13

Location: Site C0002
Weather: Cloudy

Finally BOP/Riser into the water and running smoothly to the well head.

Today's photo:
running BOP into the water.

10.18, 2018
Day:12

Location: Site C0002
Weather: Rain

BOP trouble shooting still ongoing, but seems like it ends soon. Continue reviewing lab layout and preparation for cuttings handling tools and bottles. Conduct XCT scanning and high pressure test on BOP pressure switches.

Today's photo:
Meanwhile France 5 crew are onboard for a documentary video shooting.

10.17, 2018
Day:11

Location: Site C0002
Weather: Cloudy

BOP troubleshooting continues, unfortunately. it is not a serious one but seems like takes time. Fingers crossed.
Meanwhile, preparation for cuttings handling is ongoing in the lab based on the sample request.

Today's photo:
riser pipe wating for picking up on riser transfer cart.

10.16, 2018
Day:10

Location: Site C0002
Weather: Overcast

Continue troubleshooting for BOP, run BOP/riser to 60 m water depth, recover BOP/riser to surface, troubleshooting for BOP.

Today's photo:
working on "cherry picker" above BOP

10.15, 2018
Day:9

Location: Site C0002
Weather: Fine but cloudy

Continued preparation for cuttings handling tools and bottles.

Today's photo:
the tool for wellhead cleaning we used

10.14, 2018
Day:8

Location: Site C0002
Weather: Cloudy



Dive ROV and deploy 4 transponders by 1130h.
Seabed survey around 100 m from the well center and recover ROV by 1745h Conduct field arrival check and calibration for the acoustic positioning reference system by 2045h. Meanwhile transfer BOP on BOP stand, connect Instrument Joint to Lower Marine Rise Package, re-arrange MUX cables for running BOP, and secure BOP with Moon Pool taggers. Dive ROV, clean well head and recover ROV by 2400h.

Today's photo:
well head cleaning

10.13, 2018
Day:7

Location: Site C0002
Weather: Overcast

After finish SD-RCB and RCB coring test successfully, Chikyu moved to C0002 site.After DP mode had set, ROV dived to deploy transponders on the seafloor. Cuttings treatment preparation is ongoing in the laboratory.

Today's photo:
ROV is diving into the water with transponders

10.12, 2018
Day:6

Location: SD-RCB test site C9034E
Weather: Overcast

Continuously conducts coring system (SD-RCB and RCB) in the same area. Once the test will finish, Chikyu heading to the site, C0002 immediately.

10.11, 2018
Day:5

Location: SD-RCB test site C9034D
Weather: Cloudy

The first 8-1/2” SD-RCB test core was recovered in the morning. Core was taken from the ~50mbsf, where seafloor is ~1700 m deep from the sea surface, at a site located south of the Suruga Trough, i.e. on the Philippine Sea plate.

Today's photo:
checking core recovery and appearance at Core Cutting Area.

10.10, 2018
Day:4

Location: SD-RCB test site C9034D
Weather: Rain

D/V “Chikyu” left the Shimizu Port at 10:00 am, where many people saw her off. The ship arrivedthe site for SD-RCB (small diameter-roraty coring barrel) coring test operation before move to the C0002 site in the evening. Second pre-spud meeting was held in the afternoon for the people who could not attend the first meeting.

Today's photo:
crews responded by waving hands to a grand send-off.

10.9, 2018
Day:3

Location: Shimizu Port
Weather: Fine but cloudy

Pre-spud meeting, which a Science Leader Kyu Kanagawa and key personnel of CDEX, MWJ and MQJ attended, was held in the afternoon, where information on HSE, drilling operations and scientific objectives was exchanged after a brief introduction of attendees. CDEX director, Shin’ichi Kuramoto, gave final words and a cheer.

Today's photo:
OSI Tomo Sasuhashi presented operation overview at the pre-spud meeting.

10.8, 2018
Day:2

Location: Shimizu Port
Weather: Fine but cloudy

Hold a meeting with all shipboard lab tech for cutting flow preparation.

Today's photo:
lab tech meeting in Core Cutting Area for cuttings flow preparation

10.7, 2018
Day:1

Location: Shimizu Port
Weather: Fine but cloudy

Embarked one of science leaders Kyu Kanagawa.

Today's photo:
moving casing pipes using crane